If the molarity of an hydrous sodium triox carbonate is 2.5M calculate the Mass of salt that dissolve 10cm^3 of water
1
Expert's answer
2021-08-20T01:52:40-0400

"Molarity= \frac{Moles}{Volume}"


Moles "=\frac{2.5×10}{1000}=0.025moles"


"Moles=\frac{mass}{molar-mass}"


Mass "=251.936\u00d70.025=6.3grams"